OpenStudy (anonymous):

Find the derivative of the given function. Use all lower case letters. w(q) = 9a(b^8)q The only reason I'm putting parenthesis in this problem is just to emphasize that q is not part of b^8. I'm not really sure how to do this just because there are so many letters.

OpenStudy (mathmale):

First, I appreciate your using parentheses for clarity. Good job! w(q) = 9a(b^8)q appears to be a function of q alone; I'd treat a and b as constants, at least initially.

OpenStudy (mathmale):

w(q) = 9a(b^8)q\[w(q) = 9a(b^8)q\] Is this what you meant? If not, please vix it.

OpenStudy (mathmale):

fix it, I meant.

OpenStudy (anonymous):

On Webstudy it looks like this: \[w(q)=9ab^8q\] The way it looks is what is confusing me. I can't tell if a and b are together or separate variables but just thrown together.

OpenStudy (mathmale):

Look at the label: f(q). To me, that indicates that f is a function of q alone, and that a and b are thus to be treated as constants. What would be df/dq if you had f(q)=q?

OpenStudy (anonymous):

The answer would be 1 in that case, right?

OpenStudy (mathmale):

Yes. What would be the derivative with respect to q of p(q)=(a^3)q?

OpenStudy (anonymous):

Wouldn't that be 0? Since you are multiplying q by the derivative of a constant, which is 0.

OpenStudy (mathmale):

No. Instead, you must treat (a^3) as the whole thing were a constant coefficient. Try again. What is the derivative, with respect to q, of my sample function?

OpenStudy (mathmale):

If y=5a*q, dy/dq=5a*1=5a

OpenStudy (anonymous):

Ohhh... Alright. I think I get it.

OpenStudy (mathmale):

Happy for you. Good luck! Contact me again when you have more questions.